OpenStudy (anonymous):

The sum of the angles in a convex 10-sided polygon is: a.1800° b.188° c. 3600° d. 1440°

OpenStudy (ibbutibbu.):

The correct answer is D

OpenStudy (bibby):

Why is D correct?

OpenStudy (ibbutibbu.):

The sum of the interior angles of a polygon with n sides is 2n - 4 right-angles. With n = 10, this gives: 90(2*10 - 4) = 1440deg. That makes each angle 1440/10 = 144deg.

OpenStudy (bibby):

I just did (n-2)*180 8*180=1440

OpenStudy (ibbutibbu.):

I just did it another way. Whats the diff?

OpenStudy (bibby):

The diff is that you're not supposed to give the answers. Even in multiple choice questions. You're here to explain, not do their homework

OpenStudy (ibbutibbu.):

Oh ok, im sorry. Should I delete that post?

OpenStudy (bibby):

nah it's fine, you ended up explaining

OpenStudy (ibbutibbu.):

Ok